OpenStudy (anonymous):

Find the present value of the annuity if the withdrawal is to be $600 per month for 21 months at 5% compounded monthly.

jimthompson5910 (jim_thompson5910):

\[PV = P[\frac{1-(1+i)^{-n}}{i}]\] P=600, i=0.05, n=21 \[PV =600[\frac{1-(1+0.05)^{-21}}{0.05}]\] \[PV =600[\frac{1-(1.05)^{-21}}{0.05}]\] \[PV =600[\frac{1-0.3589423646}{0.05}]\] \[PV =600[\frac{0.6410576354}{0.05}]\] \[PV =600[12.821152708]\] \[PV =7692.6916248\] \[PV =7692.69\] So the present value of the annuity is $7,692.69

OpenStudy (anonymous):

hmmmm i submitted that answer and they told me it was incorrect....

jimthompson5910 (jim_thompson5910):

k let me double check

jimthompson5910 (jim_thompson5910):

is that rate 5% the annual rate?

OpenStudy (anonymous):

im assuming its monthly since thats what the question is asking but i'm not positive

jimthompson5910 (jim_thompson5910):

how many tries you have left?

OpenStudy (anonymous):

two more! but now the values have changed

OpenStudy (anonymous):

Find the present value of the annuity if the withdrawal is to be $300 per month for 36 months at 10% compounded monthly.

jimthompson5910 (jim_thompson5910):

try 9297.37

jimthompson5910 (jim_thompson5910):

I'm sure it has to do with the interest rate

OpenStudy (anonymous):

is this for the second set of values or the first question i asked?

jimthompson5910 (jim_thompson5910):

second set

OpenStudy (anonymous):

that was correct! how did you get that??

OpenStudy (anonymous):

nevermind i got the formula! thanks so much!

jimthompson5910 (jim_thompson5910):

\[PV = P[\frac{1-(1+i)^{-n}}{i}]\] P=600, i=0.05, n=21 \[PV =300[\frac{1-(1+\frac{0.10}{12})^{-36}}{\frac{0.10}{12}}]\] \[PV =300[\frac{1-(1+0.00833333333333)^{-36}}{0.00833333333333}]\] \[PV =300[\frac{1-(1.00833333333333)^{-36}}{0.00833333333333}]\] \[PV =300[\frac{1-0.74173970345571}{0.00833333333333}]\] \[PV =300[\frac{0.25826029654429}{0.00833333333333}]\] \[PV =300[30.9912355853271]\] \[PV = 9297.37067559812\] \[PV = 9297.37\]

jimthompson5910 (jim_thompson5910):

oops I meant to write P=300, i=0.1/12, and n=36